QUESTION 8

A plane is flying at 400 mph with a heading of S 20 w. A strong cross wind is blowing at 40 mph with a heading of N 55º w. Show all steps.

a) Sketch the graph, determine the vector representing the velocity of the plane relative to the air in the form <x-component, y-component>.

b) Sketch the graph, determine the vector representing the velocity of the wind in the form <x-component, y-component>.

C) Determine the true velocity of the plane in the form <x-component, y-component>.

d) Determine the speed of the plane relative to the ground

d) What is the true bearing that the plain is flying?   
Answer #1

Page 1 400 mph Velouty of plane Velocity of wind i 520°W 40mph N55°W 200 VPG ..JPG = Velocity of plane wat ground in the formpage 2 @ VpA = Velocity g plane wit air = VPG-VWG - 375.871-136.80m +22.942 -32.77 ja = -352.93 i - 169.540j tanx=(169.57 135Page 3 True velocity = - 375.872 -136.8 6.80 tomd=1-136180 -375.87 -375.87 Q=19.8° - 136.80 (d) speed of plane relative to gr